Ask your own question, for FREE!
Mathematics 14 Online
OpenStudy (anonymous):

Jennifer flipped a pair of coins 60 times. On 24 occasions, heads appeared on both coins. How is this outcome different from the expected outcome?

OpenStudy (kropot72):

The probability of getting a head on one flip of a fair coin is 1/2. The probability of getting heads with a pair of fair coins is 1/2 * 1/2 = 1/4. The expected number of occasions of getting heads on both coins when flipping the pair 60 times is 60 * (1/4) = 15 which is much lower than the experimental number of 24. There is a possibility that one or both coins are biased.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!